# Advances in Kidney Biopsy Lesion Assessment through Dense Instance Segmentation

### v2

@@ 4.3 Complexity Analysis
We start with analyzing the class imbalance within our dataset. Of the total structures ($21,889$ objects), 88% are tubuli, followed by 7% arteries (1679 objects), and 5% glomeruli (1226 objects). This distribution is visualized in Figure.11(a), highlighting the necessity of sampling to ensure the inclusion of rare instances, such as arteries and glomeruli, among the dominant tubuli. A violin plot, shown in Figure.11(b), details the distribution of structure counts at the ROI level. On average, ROIs contain $95$ objects, with a maximum of $434$. To accommodate all objects, DiffRegFormer is configured to handle a maximum number exceeding $434$ per image. The correlation between the increase in computational cost and performance improvement, depicted in Figure 11(c). As expected, we observe a turning point at $500$ dynamic queries. This configuration yields a favorable trade-off between complexity and efficiency, achieving 52.1% AP with $20.2$GB GPU memory usage. Specifically, ROIs contain an average of $5$ glomeruli, $7$ arteries, and $84$ tubuli with corresponding maximums of $74$, $84$, and $408$. Setting the maximum number per class to $133$ is another feasible option for positive sample generation, as our sampling approach ensures comprehensive inclusion of rare instances while maintaining rational proportion of the majority class. In conclusion, selecting a maximum of $500$ objects per image and $133$ per class achieves an advantageous balance between performance and computational efficiency.
@@ 4.4 Ablation Study
In this ablation study, we focus on the effects of various training strategy combinations and query types on model performance. Through analysis of these results, we can gain insights into the specific function of each strategy. This analysis highlights the significance of our improvements as key technical contributions for adapting an end-to-end instance segmentation model to datasets with dense objects.
@@ 4.4.1 Training strategies
Training an instance segmentation model for datasets with dense objects, such as kidney biopsies, presents distinct challenges compared to datasets with sparser object distributions like MSCOCO. Instance segmentation models typically employ three strategies: (1) shared features between the bbox-decoder and mask-decoder; (2) selection of positive samples from proposal bounding boxes; and (3) class-imbalanced sampling. However, for dense object segmentation, we propose three alternative strategies: (a) separate features for the bbox and mask decoders; (b) positive sample selection directly from ground truth boxes; and (c) class-balanced sampling. While the conventional combination of shared features, proposal boxes, and biased sampling is effective for standard instance segmentation (Figure.12a), only the combination of separate features, ground-truth boxes, and unbiased sampling proves successful for dense instance segmentation. Table.4 presents an evaluation of these combinations, with conventional strategies marked with ✗ and our proposed strategies with ✓. Notably, only five out of eight possible combinations can train DiffRegFormer.
@@
Figure.12b demonstrates that the combination of shared features, ground-truth boxes, and biased sampling enables object detection across all classes. However, it fails to produce accurate instance masks within these bounding boxes. This suggests that the bbox decoder dominates the shared feature representation, hindering the mask decoder’s ability to extract pixel-wise information. Consequently, all pixels within detected boxes are classified as tubuli. Moreover, biased sampling impedes the mask decoder’s capacity to learn features for rare instances (glomeruli and arteries), resulting in no pixel classification within those boxes.
@@
Additionally, Figure.12c, shared features combined with ground-truth boxes and unbiased sampling can detect and segment objects across all classes, albeit with the degraded mask quality. This degradation is likely due to the bbox decoder’s dominance within the shared feature representation.
@@
In contrast, Figure.12d reveals that using separate features, proposal boxes, and unbiased sampling leads to inaccurate object detection. This is likely due to the accumulation of errors introduced by proposals during the early stages of diffusion model training.
@@
Lastly, Figure.12e demonstrates that separate features, ground-truth boxes, and biased sampling result in correct detection and segmentation for only a small proportion of tubuli. This finding further highlights the challenges posed by biased sampling in learning features for rare instances.
@@
In summary, shared feature maps lead to the bbox-decoder dominating feature representation learning, impairing the mask-decoder’s ability to extract pixel-wise information necessary for accurate binary instance masks. Biased sampling further hinders the mask-decoder to learn information for segmenting rare instances (glomeruli and arteries), causing it to be overwhelmed by the majority class (tubuli). Finally, utilizing proposals for sampling instead of ground truth boxes introduces cumulative errors that disrupt the training for both decoders, preventing to learn contextual information necessary for long-range correlations via attention mechanisms.
@@ 4.4.2 Queries
Global queries in transformer-based models face challenges for representing dense objects within large-scale datasets due to their static nature. Effectively accommodation to extensive variations in object size and shape necessitates a proportional increase in query count. Conversely, dynamic queries, generated on-the-fly for each object in a mini-batch, demonstrate superior adaptability to dataset complexity. Theoretically, assigning a larger number of global queries could achieve comparable performance to dynamic queries. However, our experiments demonstrate that increasing the number of global queries without a corresponding growth in feature channels does not yield improvements. Consequently, employing global queries for dense object processing in transformer-based models is significantly less cost-effective than utilizing dynamic queries.
Table.5 delineates the comparative analysis of varying query counts for both dynamic and global queries, with the number of feature channels fixed at $256$. We observe that increasing the number of dynamic queries leads to performance gains, while increasing the number of global queries results in performance degradation. Figure.13 further illustrates the inherent challenge of achieving a favorable trade-off between complexity and performance with global queries. Even with $500$ queries, the model with global queries fails to consistently detect structures across all classes, a limitation not attributable to class imbalance, in contrast to the dynamic query strategy. In conclusion, using dynamic query is a more efficient approach for datasets with dense objects, highlighting their superiority over global queries.
@@ 5 Conclusion and Future Work
This research introduces a novel framework to process large-scale datasets with potential changes in lesion combinations. Our model demonstrates superior performance in predicting lesions within dense objects, effectively handling multi-class and multi-scale challenges at the ROI level. We present the first dense instance segmentation module, DiffRegFormer, that seamlessly integrates a diffusion model with a transformer in RCNN style. This approach leverages the diffusion model to generate refined object boxes, upon which a regional transformer creates dynamic queries that are ultimately transformed into accurate instance masks within the boxes. Furthermore, our lesion classifier accurately predicts class-specific lesions for each cropped structure. The model demonstrates several advantages in handling dense object datasets: (1) its iterative noise-to-box approach eliminates reliance on prior knowledge about object size and shape, facilitating rapid object localization; (2) dynamic queries, generated on-the-fly for each object in a mini-batch, ensures scalability to large datasets; (3) the model’s modular design, with independent plug-in prediction heads, allows for replacement and promotes adaptability. This flexibility enables efficient reuse of trained components when scaling to large-scale datasets. Experimental results on Jones’ stained renal biopsies demonstrate that our model surpasses existing benchmarks.
While the current model may not yet meet the rigorous requirements for clinical applications, it establishes a robust framework for dense, multi-class, multi-scale object recognition at the ROI level, setting a solid foundation for future engineering improvements. Potential areas for future research include: (1) further enhancing the compatibility of diffusion models with transformers. (2) developing innovative regional feature extractors to generate more informative dynamic queries and mitigate the loss of information for small objects during iterative refinements. (3) integrating spatial constraints into regional mask inference to prevent erroneous assignment of single pixels to multiple instances.

### v3

@@ 4.3 Complexity Assessment of Model Training
We start with analyzing the intrinsic class imbalance within our dataset. Of the total structures ($21,889$ objects), 88% are tubuli, followed by 7% arteries (1679 objects), and 5% glomeruli (1226 objects). This distribution is depicted in Figure 11a, highlighting the necessity of sampling to include more rare instances, such as arteries and glomeruli, among the dominant tubuli. A violin plot, shown in Figure 11b, details the distribution of structure counts at the ROI level. On average, ROIs contain $95$ objects, with a maximum of $434$. To accommodate all objects, DiffRegFormer is configured to handle a maximum number exceeding $434$ per image. Figure 11c depicts the correlation between the increase in computational cost and performance improvement. As expected, we observe a turning point at $500$ dynamic queries. This configuration yields a favorable trade-off between complexity and efficiency, achieving 52.1% AP with $20.2$GB GPU memory usage. Specifically, ROIs contain an average of $5$ glomeruli, $7$ arteries, and $84$ tubuli with corresponding maximums of $74$, $84$, and $408$. Setting the maximum number per class to $166$ is another feasible option for positive sample generation, as our sampling approach ensures comprehensive inclusion of rare instances while maintaining the rational proportion of the majority class. In conclusion, selecting a maximum of $500$ objects per image and $166$ per class can effectively balance performance and computational efficiency.
@@ 4.4 Ablation Study
The ablation study focuses on the effects of various training strategies and query types on model performance. By analyzing our results, we can gain insights into the specific characteristics of each strategy. This analysis highlights the significance of our improvements as key contributions to adapting an end-to-end instance segmentation model to datasets with dense objects.
@@ 4.4.1 Training strategies
Training an instance segmentation model for datasets with dense objects, such as, in our case, kidney biopsies, presents distinct challenges compared to the common datasets with sparser object distributions like MSCOCO. Instance segmentation models typically employ three strategies: (1) shared features between the bbox-decoder and mask-decoder; (2) selection of positive samples from proposal bounding boxes; and (3) class-imbalanced sampling. However, for dense object segmentation, we propose three alternative strategies: (a) separate features for the bbox and mask decoders; (b) positive sample selection directly from ground-truth boxes; and (c) class-balanced sampling. While the conventional combination of shared features, proposal boxes, and biased sampling is adequate for standard instance segmentation (see Figure 12a), only the combination of separate features, ground-truth boxes, and unbiased sampling proves to be successful for dense instance segmentation. Table 4 evaluates the combinations above, with conventional strategies marked with ✗ and our proposed strategy with ✓. Notably, only five out of eight possible combinations can train DiffRegFormer.
@@
Figure 12b demonstrates that combining shared features, ground-truth boxes, and biased sampling enables object detection across all classes. However, it fails to produce accurate instance masks within these bounding boxes. The observation suggests that the bbox-decoder dominates the shared feature representation, thus hindering the mask decoder’s ability to extract pixel-wise information. Consequently, all pixels within detected boxes are classified as tubuli. Moreover, biased sampling impedes the mask decoder’s capacity to learn features for rare instances (glomeruli and arteries), resulting in no pixel classification within those boxes.
@@
Additionally, Figure 12c, shared features combined with ground-truth boxes and unbiased sampling can detect and segment objects across all classes, albeit the degraded mask quality. This degradation is likely due to the bbox-decoder’s dominance within the shared feature representation.
@@
In contrast, Figure 12d reveals that using separate features, proposal boxes, and unbiased sampling leads to inaccurate object detection. It is probably due to the accumulation of errors introduced by proposals during the early stages of diffusion model training.
@@
Lastly, Figure 12e demonstrates that separate features, ground-truth boxes, and biased sampling result in correct detection and segmentation for only a small proportion of tubuli. This finding further highlights the challenges caused by biased sampling in learning features for rare instances.
@@
In summary, shared feature maps lead to the bbox-decoder dominating feature representation learning, impairing the mask-decoder’s ability to extract pixel-wise information necessary for accurate binary instance masks. Biased sampling further hinders the mask-decoder from learning information to segment rare instances (glomeruli and arteries), causing it to be overwhelmed by the majority class (tubuli). Finally, utilizing proposals for sampling instead of ground-truth boxes introduces cumulative errors that disrupt the training for both decoders, preventing the learning of contextual information necessary for long-range correlations via attention mechanisms.
@@ 4.4.2 Queries
Static queries in transformer-based models face challenges in representing dense objects within large-scale datasets due to their static nature. Adequate accommodation to extensive variations in object size and shape necessitates a proportional increase in query count. Alternatively, dynamic queries, generated on-the-fly for each object in a mini-batch, demonstrate superior adaptability to dataset complexity. Theoretically, assigning more static queries could achieve comparable performance to dynamic queries. However, our experiments indicate that increasing the number of static queries without a corresponding increase in feature channels does not yield improvements. Consequently, employing static queries for dense object processing in transformer-based models is significantly less cost-effective than dynamic queries.
Table 5 delineates our comparative analysis of varying query counts for both dynamic and static queries, with the number of feature channels fixed at $256$. We observe that increasing the number of dynamic queries leads to a gain in performance while increasing the number of static queries results in performance degradation. Figure 13 further illustrates the inherent challenge of achieving a favorable trade-off between complexity and performance with static queries. Even with $500$ queries, the model with static queries fails to detect structures across all classes consistently. This limitation is not attributable to class imbalance compared to the dynamic query strategy. In conclusion, the dynamic query is a more efficient approach for datasets with dense objects, highlighting their superiority over static queries.
@@ 5 Conclusion and Future Work
Our research introduces a novel framework to process large-scale renal WSI datasets with potential changes in lesion combinations. Our model performs better in predicting lesions within dense objects, effectively handling multi-class and multi-scale challenges at the ROI level. We present the first dense instance segmentation module, DiffRegFormer, that seamlessly integrates a diffusion model with a transformer in RCNN-style. This approach leverages the diffusion model to generate refined object boxes, upon which a regional transformer creates dynamic queries that are ultimately transformed into accurate instance masks within the boxes. Furthermore, our lesion classifier accurately predicts class-specific lesions for each cropped structure. The model demonstrates several advantages in handling dense object datasets: (1) its iterative noise-to-box approach eliminates the dependence on prior knowledge about object size and shape, facilitating rapid object localization; (2) dynamic queries, generated on-the-fly for each object in a mini-batch style, ensuring scalability to large datasets; (3) the modular design of the model, with independent plug-in prediction heads, allows for replacement and promotes adaptability. This flexibility enables the efficient reuse of trained components when up-scaling to large-scale datasets. Experimental results on Jones’ stained renal biopsies demonstrate that our model surpasses existing benchmarks.
The current model needs further exploration to meet the rigorous requirements for clinical applications. Still, it establishes a robust framework for dense, multi-class, multi-scale object recognition at the ROI level, setting a solid foundation for future computational improvements. Potential areas for future research include: (1) further enhancing the compatibility of diffusion models with transformers like DETR to avoid NMS post-processing. (2) developing innovative regional feature extractors to generate more informative dynamic queries and mitigate the loss of information for small objects during iterative refinements. (3) integrating spatial constraints into regional mask inference to prevent erroneous assignment of single pixels to multiple instances. (4) designing lightweight attention mechanisms to lessen the computational burden. In conclusion, the approach presented paves the way for efficient and reliable automated assessment as a tool for the workbench of renal pathologists.
@@ 6 Acknowledgements
Appendix A Additional Visualization Samples
As an extra, we have added two zoom-in samples to highlight the details of our method and other approaches. The first row of Figure A.17 shows that our method can correctly divide two glomeruli without overlap. Cascade Mask RCNN and Mask RCNN, however, generate multiple responses at the border of two objects. QueryInst with static queries cannot detect and segment any glomerulus. The last row of Figure A.17 demonstrates four large arteries in a line. It is complicated since two large arteries pass over most input image regions. Although it had slightly multiple responses at the artery border, our method can capture long-range spatial dependencies for large objects using attention mechanisms with dynamic queries. In contrast, Cascade Mask RCNN and Mask RCNN cannot model large objects without attention modules. QueryInst adopts an attention mechanism with static queries and cannot process complicated combinations of shapes and appearances with fixed query numbers.
Fig. 17: Detailed zoom-in samples comparing DiffRegFormer with other methods.